Grade 12th PassAlgebra

Find all the values of the parameter a for which the inequality 4x - a2x -a+3≤0 is satisfied by at least one real x.

Menka Malguri
14 Years agoGrade 12th Pass

1 Answer

Chetan Mandayam Nayakar
ApprovedApproved Tutor Answer14 Years ago

let 2x=y, y is always non-negative, a≥(y^2 +3)/(y+1) =f(y), f(y) is minimum when 2y(y+1)=y^2 +3, y=1 (y=-3 is ruled out because y=2^x is never negative for real x), f(1)=2, therefore answer is a≥2